OpenStudy (anonymous):

How do I solve this equation: ln(x+1)/ln(x-1) = 2?

OpenStudy (anonymous):

that equation can be written as: \(\large ln(x+1)=2ln(x-1) \) or \(\large ln(x+1)=ln(x-1)^2 \rightarrow (x+1)=(x-1)^2\) agreed?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so can you solve that last equation: \(\large (x+1)=(x-1)^2 \) ???